### v1

# On the Readiness of Scientific Data for a Fair and Transparent Use in Machine Learning

Abstract
To ensure the fairness and trustworthiness of machine learning (ML) systems, recent legislative initiatives and relevant research in the ML community have pointed out the need to document the data used to train ML models. Besides, data-sharing practices in many scientific domains have evolved in recent years for reproducibility purposes. In this sense, the adoption of these practices by academic institutions has encouraged researchers to publish their data and technical documentation in peer-reviewed publications such as data papers. In this study, we analyze how this scientific data documentation meets the needs of the ML community and regulatory bodies for its use in ML technologies. We examine a sample of 4041 data papers of different domains, assessing their completeness and coverage of the requested dimensions, and trends in recent years, putting special emphasis on the most and least documented dimensions. As a result, we propose a set of recommendation guidelines for data creators and scientific data publishers to increase their data’s preparedness for its transparent and fairer use in ML technologies.
1 Introduction
The growing influence of machine learning (ML) technologies in our society raises scientific and political concerns about the potential harms they may cause. Datasets are critical in these systems, and recent research has highlighted them as one of the root causes of unexpected and harmful consequences in ML applications. Recent studies, for example, observed gender-biased classifiers for computer-aided diagnosis due to imbalances in training data [1, 2], or ML models for pneumonia detection failing to generalize to other hospitals due to specific conditions during the collection of the images used to train them [3]. This situation has prompted the interest of regulatory agencies and the ML community in general in developing data best practices, such as building proper dataset documentation. Public regulatory initiatives, such as the European AI ActaaaAI Act required documentation: Annex IV: https://www.euaiact.com/annex/4 and the AI Right of Billsbbbhttps://www.whitehouse.gov/ostp/ai-bill-of-rights/, as well as relevant scientific works [4, 5], have proposed general guidelines for developing standard dataset documentation. Such proposals identify a number of dimensions, such as the dataset’s provenance or potential social issues, that could influence how the dataset is used and the quality and generalization of the ML models trained with it.
Interest in data-sharing practices was already widespread in many scientific fields, determined, among others, by the need to reproduce scientific experiments [6]. The adoption of FAIR principles [7] and data management plans [8] by research institutions has encouraged researchers to publish their data with technical and scientific documentation, such as data papers. Recent research has focused on how to improve data-sharing practices through this type of documents, working on specific aspects such as the peer-review process for data papers [9], data citation practices [10], data publishers’ guidelines [11], and data reuse [12].
The use of this type of data for the training of ML models is making good data practices even more important. However, research has yet to be conducted on the suitability of these data-sharing practices and the data dimensions required by the ML community. In this sense, the goal of this paper is to study the presence (and evolution) of such dimensions in current data papers published by the scientific community. To do so, we analyzed 4041 open access data papers published in two interdisciplinary data journals: Nature’s Scientific Data and Elsevier’s Data in Brief. We examined the full manuscripts regarding the demanded dimensions which are detailed in Section 2. The analysis was aided by an ML pipeline that included a Large Language Model (LLM) at its core, designed to extract specific dimensions from the manuscripts.
As we detail in our results—cf. Section 3—we noticed a significant contrast in the presence of the required dimensions. Those already explicitly requested in the publisher submission guidelines, such as the data recommended uses or a general description of the collection process, are consistently present. In contrast, details about the data generalization limits and its social concerns, the profiles of the team collecting and annotating the data, and the data maintenance policies need to be much better documented if we want to ensure ML models trained on trustworthy and reliable data. To help with this, in Section 4, we propose a set of recommendation guidelines for data creators and scientific data publishers to increase the presence of the least documented dimensions as a way to improve their data for its transparent and fair use in machine learning. The ML extraction pipeline and the entire dataset used in the analysis are openly available [13].
2 Background
The need for proper documentation for datasets in the ML field is well defined in the well-known publication Datasheets for Datasets[4], a work inspired by datasheets in the electronics industry. In this work (among others), the authors identify data aspects for each phase of the data creation process—design, gathering, and annotations—that could affect how the dataset should be used or the quality of ML models trained. In Table 1, we summarize the main dimensions covered in this section.
Table 1: Dimensions analyzed from the sample of scientific data papers
Dimensions Subdimensions Target explanation
Uses Recommended uses Recommended uses and gaps the dataset intends to fill
Generalization limits Non-recommended uses and data generalization limits
Social concerns If represent people: biases, sensitivity and privacy issues
Maintenance Policies Maintainers & policies (erratum, updates, deprecation)
Tested using an ML approach Models and metrics the dataset have been tested on
Collection Description Description of the process and its categorization
Profile of the collection team Profile of the gathering team
Profile of the collection target If represent people, their demographic profile
Speech context in language datasets If language datasets, contextual information of the spoken language
Collection sources & infrastructure The source of the data and the infrastructure used to collect it
Annotation Description Description of the process and its categorization
Profile of the annotator team Profile of annotation the team
Annotation infrastructure The tools used to annotate the data
Annotation validation Validation methods applied over the annotations
The Uses dimensions of the data refer to how this data should or should not be used and its potential social effects. In terms of Recommended uses, Datasheets for Datasets, among other works[14, 15, 16], propose to document the purposes and gaps the dataset intends to fill. Also, they demand rationales about the Generalization limits of the data and the potential Social concerns if the data represents people. For instance, data gathered from patients of a single hospital may have generalization limits for various reasons. The most obvious is data distribution, where, for instance, a cohort of patients is not representative of the whole population [3]. Also, some of the ML community works [17] express concerns about the Maintenance of the data. Setting an erratum to inform potential errors, an updating timeframe, or deprecation policies are remarked as useful practices while releasing data.
Data provenance is also the focus of these works. Documenting specific details about the Collection and Annotation process is stated as a need to evaluate properly the suitability of the data for specific applications. For instance, recent studies [3] have shown that the use of specific devices and practices during the collection of chest lung cancer images could raise a biases in data, leading to spurious correlations [18] of the ML model, and providing incorrect accuracy metrics. These examples illustrate the usefulness of knowing information about the Sources & Infrastructure, such as the scanners used to get MRI images, and other contextual attributes of the data. As another example, the annotation Infrastructure used, such as a labeling software, and the Validation methods applied over the labels could reduce the potential bias of the data and could give us clues about the data generalization limits. In that sense, a data user may provide more confidence to a record of data labeled by a group of people together with the inter-annotator agreement between them than labeled by a single one.
Also, there is a demand to characterize the people who have participated in the provenance steps of the data, as they add subjectivity to it [19]. For instance, a dataset curated by a team of experts in the field or curated by a crowdsourcing service, such as Amazon Mechanical Turk, could not assign the same values. Therefore standalone data without information about the team involved in its creation may not be useful to evaluate the suitability of the data for an application. As creating data involves more people every time, there is a need to go beyond “the authors” of the dataset and Profile the collection team and the annotation team of the data. Beyond those who actively created the data, recent works [20] also stress the need to profile passive participants of the dataset creation. For instance, if the data is collected by or represents people, a profile of these people should also be required (Profile of the collection target). This practise is widely integrated in health data where the profile of the cohort of patients is highly relevant. Finally, for natural language data, which has gained relevance to the recent emergence of large language models, recent works [15] also stress the need to document contextual information about the speech (Speech Context) such as the specific dialect or the modality of the text. For instance, a natural language dataset gathered from Australian speakers may reduce the accuracy of models trained to support users in the United States due to different accents and language communication styles.
3 Results
This section reports on the results of our analysis. We first provide some information on the representativeness of the sample, describing the size and diversity of the data papers, and second, we report on the presence or absence of the information dimensions that characterize a good dataset for ML training purposes. We also provide a temporal evolution of such data to see whether the situation is improving.
3.1 Size and diversity of the sample
The analyzed sample comprises 4041 data papers extracted from two interdisciplinary scientific journals. In Figure 1, we can see an apparent growing trend in the number of papers published between 2015 and 2023, with 2021 being an exception due to the effect of COVID-19 on general scientific production. Note also that the absolute value for 2003 is smaller as we only collected data until June of 2023. The 16,5% of the sample’s data papers are related to people, and this is the subsample we used to calculate the presence of the Social concerns and the Profile of the collection targets dimensions in the next section as these metrics are only relevant for people-related datasets.
As an indication of the diversity of our dataset, Table 3.1 shows the 10 most popular topics of the uses of the data in the analyzed sample, showing the sample diversity. The topics related to the Conservation of biodiversity, RNA sequencing for gene analysis, and Chemistry and Physics are the most common ones. Other topics, such as Breast Cancer and Magnetic Resonance Imaging (MRI), are very interrelated, where most of the datasets around Breast Cancer are composed of MRI images. On the other hand, topics such as COVID-19 have appeared quickly in 2021 and are decreasing over time, as they are strongly correlated with a temporal event such as the coronavirus pandemic. The full results of the topic analysis can be found in our open repository [13].
Figure 1: Number of data papers published between 2015 and 2023 evaluated in the sample. 2023 has been evaluated until June.
Table 2: Uses: Extract of the 10 most popular topics in data uses of the analyzed data papers
Topic
Related data papers (Sample’s percentage)
Example
Conservation of biodiversity 352 (8,7%)
"To study the biodiversity and distribution of fish species, and assess the conservation status of fish species." [21]
RNA sequencing for gene analysis 297 (7,4%)
"Quality control and data analysis in RNA-seq experiments, including quality analysis of raw sequencing data…" [22]
Chemistry and physics 203 (5,0%)
"Soil chemistry analysis, soil classification, and land use analysis in New Zealand’s major agricultural regions." [23]
Breast Cancer 181 (4,5%)
"MRI image processing procedures and constructing breast models for studying the effects of chemotherapy on breast cancer patients." [24]
Magnetic resonance imaging 142 (3,5%)
"Preprocessing of functional MRI data, including motion correction, slice-timing correction, non-brain removal…" [25]
Covid19 104 (2,6%)
"To identify geographies for surveillance and inform global health security…" [26]
Smart Homes 61 (1,5%)
"Gain a deep understanding of residential energy use, analyze household behavior and opinions on energy choices…" [27]
Precipitation 60 (1,5%)
"Comparing different precipitation datasets with observations from GHCN-D 80 for the continental United States except Alaska…" [28]
Data Urban Transport 56 (1,4%)
"Analyzing the impacts of congestion in functional urban areas, benchmarking functional urban areas according to accessibility…" [29]
Cropland 54 (1,3%)
"Optimizing agricultural production strategies and water resources management, as well as validating crop yield…" [30]
Figure 2: Diversity in the collection and annotation teams type of the analyzed data papers
Figure 3: Collection: Diversity of types of collection processes
Figure 2: Diversity in the collection and annotation teams type of the analyzed data papers
Figure 3: Collection: Diversity of types of collection processes
Figure 4: Annotation: Diversity of types of annotation processes
Our sample shows diversity in terms of the annotation and collection processes types as well. For instance, Figure 4 depicts the distribution of the most common data collection types with Physical data collection, Direct measurements, Document analysis, Manual human curator, and Software collection (all above 10%) being the most common ones. Moreover, Table 4 shows that nearly 88,05% of the teams who participated in data collection were internal teams (authors of the study or close collaborators), 11,20% were external teams, and only 0,74% used crowdsourcing practices to collect the data. Regarding the annotation process, only 42,28% of the analyzed data papers employ an annotation process involving humans. Figure 4 shows that the most common type of annotation is Text Entity Annotation, followed by Semantic segmentation, Polygonal segmentation and Bounding boxes. Compared to the collection processes, 29% of the processes were classified as Others. The results for team type are similar to the results for the collection process, with a high percentage of external teams (11,27%) and crowdsourcing teams (1,90%).